OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (anonymous):

@WordGEEK

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

Ok you want to isolate x in the first equation. You have -x-2y=4 Divide both sides by -1 so that you can get x to be positive. You would then have x+2y=4 Subtract 2y from both sides and you will have isolated x which results in x=4-2y or in other words answer choice 4. -2y+4